# Digital Simulation of Non-Abelian Parafermions in Superconducting Circuits

### v4

@@ VI Experimental parameters
In this section, we briefly discuss typical parameters for SC circuits to realize our scheme, particularly for the case of $d=3$. One key parameter in SC circuits is the fidelity of single- and two-qutrit gates. Recent work has demonstrated the realization of the generalized Hadamard gate for qutrits, achieving a high fidelity of $0.992$ [82]. More challenging is the implementation of high-fidelity two-qutrit gates. In recent study Ref. [84], the authors report the realization of $CZ^{\dagger}$ and $CZ$ gates for qutrits, with fidelities of $0.973$ and $0.952$, respectively. Since the fidelity of single-qutrit gates is much higher than that of two-qutrit gates, we will, as a first approximation, neglect errors from single-qutrit gates in the following analysis. Our ground state preparation protocol requires three two-qutrit gates for each dark plaquette. To achieve a ground state preparation fidelity of $0.90$, the fidelity of the two-qutrit gates needs to be above $0.965$, which is very close to the fidelity achievable with existing technology. Achieving higher level of fidelity would require further technological advances in SC devices, highlighting an important avenue for future research in this field. Other typical parameters include the relaxation time $T^{ij}_{1}$ and the dephasing time $T^{ij}_{2}$ between different levels ($ij\in\{01,12,02\}$), which must be sufficiently long to complete the braiding, fusion procedure, and measurement. Recent studies on transmon qutrits report $T^{ij}_{1}$ and $T^{ij}_{2}$ values typically around $10\mu s$ or higher [81, 84]. Considering that the gate times for high precision control are on the order of $10-100ns$ for single-qutrit gates and $100-800ns$ for two-qutrit gates, these parameters suggest that ground state preparation, along with subsequent fusion and braiding procedures, is highly promising and achievable with the near-future technology [81].
In the main text, we consider a $6\times 6$ qutrit system for illustration, though this is not the minimal setup for our scheme. To stduy the fusion rules, a $3\times 4$ qutrit setup is sufficient, while for braiding, a $4\times 4$ setup is adequate. Reducing the number of qutrits in each row (or column) decreases the number of parafermion movement operations by two, saving the operation time by hundreds to thousands of nanoseconds. This reduction significantly enhances the success rate and fidelity of the experiment. The $6\times 6$ example allows for quantum information processing based on more parafermions, such as $8$ parafermions encoding two logical qutrits and their braiding. Additionally, considering braiding on well-designed quasi-one-dimensional configurations can further minimize the required number of qutrits.
@@ VII Discussion and conclusion
Before concluding, we discuss the potential applications in high-dimensional quantum computing. Recent developments highlight several advantages over qudit-based quantum information processing [96] and flexible simulation of quantum dynamics with qudits [97]. Qudit versions of various quantum algorithms have already been developed [98, 99, 89]. High-dimensional quantum computing based on non-Abelian parafermions is inherently fault-tolerant due to topological protection. While realizing parafermions in physical materials remains a great challenge, our digital simulation scheme offers a promising approach to high-dimensional gate-based topological quantum computing in the near future. One exciting direction is to implement qudit quantum algorithms using parafermions braidings. Another challenge is using digital simulation of parafermions to implement the topologically protected $CX$ gate, involving eight parafermions and twelve half-braidings [56].
In summary, we have proposed in this work an experimental scheme for realizing and manipulating non-Abelian parafermions in SC circuits, which should be also applicable to other quantum simulation and quantum computing platforms, like trapped ions and neutral atom arrays which develop fast recently. We have demonstrated efficient protocols for ground state preparation and parafermion creation, and proposed a generalized code deformation approach to achieve the fusion rules and braiding statistics through experimental observables. This work with experimental feasibility paves the way for the digital simulation of parafermions and their braiding statistics in SC circuits.
@@ Acknowledgments
Among the excitation states, we identify that states with topological charge $\{0,1,\dots,0\}$ on a given light (dark) plaquette $p$ correspond to a charge (flux) excitation $e$ ($m$) on $p$. Additional quantum numbers $\pi_{p}^{\lambda}=1$ indicate the presence of a charge (flux) excitation $e^{\lambda}$ ($m^{\lambda}$) on the light (dark) plaquette $p$. For convenience, we define $\overline{e}\equiv e^{d-1}$ and $\overline{m}\equiv m^{d-1}$. These elementary excitations can be created in pairs and moved via string operators. More generally, composite excitations are formed from these elementary excitations.
@@
where $F$ is the Fourier transformation operator defined by $F=\frac{1}{\sqrt{3}}\sum_{k,m=0}^{2}\omega^{km}|\overline{k}\rangle\left\langle\left.\overline{l}\right|\right.$. Finally, we obtain the matrix of $U_{2}^{2}$ in Eq. (19).

### v5

@@ VI Experimental parameters
In this section, we briefly discuss typical parameters for SC circuits relevant to the implementation of our scheme, focusing on the case of $d=3$. A key parameter in SC circuits is the fidelity of single- and two-qutrit gates. Recent experiments have demonstrated the generalized Hadamard gate for qutrits with a high fidelity of $0.992$ [144]. Implementing high-fidelity two-qutrit gates is more challenging. A recent study reported fidelities of $0.973$ and $0.952$ for the realization of $CZ^{\dagger}$ and $CZ$ gates of qutrits, respectively [146]. Since single-qutrit gate fidelities are considerably higher, we neglect their contribution to overall errors in our analysis. Our ground state preparation protocol requires three two-qutrit gates per dark plaquette. To achieve a total fidelity of $0.90$ for ground state preparation, the fidelity of each two-qutrit gate needs to exceed $0.965$. This value is already within reach of current experimental platforms. Achieving even higher fidelities would require further advances in SC technology, marking an important direction for future research. Other important parameters include the relaxation time $T^{ij}_{1}$ and the dephasing time $T^{ij}_{2}$ between different qutrit levels ($ij\in\{01,12,02\}$). These times must be sufficiently long to allow for fusion, braiding, and measurement procedures. Recent experiments on transmon qutrits report typical values of both $T^{ij}_{1}$ and $T^{ij}_{2}$ around $10\mu s$ or higher [143, 146]. Considering that single-qutrit gate durations are on the order of $10-100ns$ and two-qutrit gates take about $100-800ns$, these coherence times suggest that ground state preparation and subsequent parafermion manipulations are feasible with near-term SC technologies. [143].
In the main text, we use a $6\times 6$ qutrit array for illustrative purposes, though this is not the minimal configuration required by our scheme. For studying fusion rules, a $3\times 4$ array suffices, while braiding operations can be demonstrated with a $4\times 4$ setup. Reducing the number of qutrits per row or column lowers the total number of parafermion movement operations by two, saving hundreds to thousands of nanoseconds in operation time. This reduction leads to a significant improvement in both the success rate and the overall fidelity of the experiment. The $6\times 6$ configuration, however, enables more advanced protocols, such as encoding two logical qutrits using eight parafermions and performing their braiding operations. Moreover, by designing quasi-one-dimensional layouts optimized for braiding, it is possible to further minimize the required number of qutrits.
@@ VII Discussion and conclusion
Before concluding, we discuss potential applications in high-dimensional quantum computing. Recent developments highlight several advantages over qudit-based quantum information processing [154] and flexible simulation of quantum dynamics with qudits [100]. Qudit versions of various quantum algorithms have already been developed [155, 156, 149]. High-dimensional quantum computing based on non-Abelian parafermions is inherently fault-tolerant due to topological protection. While realizing parafermions in physical materials remains a great challenge, our quantum simulation scheme offers a promising approach to high-dimensional topological quantum computing in the near future. One exciting direction is to implement qudit quantum algorithms using parafermions braidings. Another challenge is to implement the topologically protected $CX$ gate in quantum simulators, involving eight parafermions and twelve half-braidings [107].
In summary, we have proposed in this work a novel experimental scheme for realizing and manipulating non-Abelian parafermions in SC circuits, which is also applicable to other quantum simulation and quantum computing platforms, including the trapped ions and neutral atom arrays which develop fast recently. We have demonstrated efficient protocols for ground state preparation and parafermion creation, and proposed a generalized code deformation approach to achieve the fusion rules and braiding statistics through experimental observables. This work with experimental feasibility paves the way for the quantum simulation of parafermions and their non-Abelian statistics in SC circuits.
@@ Acknowledgments
Among the excitation states, we identify that states with topological charge $\{0,1,\dots,0\}$ on a given light (dark) plaquette $p$ correspond to a charge (flux) excitation $e$ ($m$) on $p$. Additional quantum numbers $\pi_{p}^{\lambda}=1$ indicate the presence of a charge (flux) excitation $e^{\lambda}$ ($m^{\lambda}$) on the light (dark) plaquette $p$. For convenience, we define $\overline{e}\equiv e^{d-1}$ and $\overline{m}\equiv m^{d-1}$. These elementary excitations can be created in pairs and moved via string operators. More general composite excitations are formed from these elementary excitations.
@@
where $F$ is the Fourier transformation operator $F=\frac{1}{\sqrt{3}}\sum_{k,m=0}^{2}\omega^{km}|\overline{k}\rangle\left\langle\left.\overline{l}\right|\right.$. Then, we obtain the full braiding matrix $U_{2}^{2}$ as shown in Eq. (19).